OpenStudy (anonymous):

A certain business has implemented a new incentive plan whereby an employee earns tokens for every sale they make. For 1 sale, an employee earns 1 token. For 2 sales, an employee earns 9 tokens. For 3 sales, an employee earns 21 tokens. For 4 sales, an employee earns 37 tokens. Which recursive equation represents the pattern? an = an - 1 + 4n an = an - 1 + 4(n + 1) an = an - 1 + (n + 4) an = an - 1 + (n – 4)

OpenStudy (anonymous):

anyone

OpenStudy (anonymous):

an = an - 1 + 4n

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

yes just replace some numbers for example to find the second term a2 u can do this a1=1 a2=1+4*2 =9 u can see that that is the correct values

OpenStudy (anonymous):

ok thanks